OpenStudy (anonymous):

Let h(x)=x^3 + 1 Determine functions f(x) and g(x) such that h(x)=f(g(x)) [Do not use f(x)=x or g(x)=x]

OpenStudy (unklerhaukus):

\[g(x)=x^3\]\[f(x)=x+1\] \[h(x)=f \circ g=f(g(x))=f(x^3)=x^3+1\]

OpenStudy (anonymous):

That makes sense, thank you for your help.

OpenStudy (unklerhaukus):

i think there could be a few different \(f(x)\)'s and \(g(x)\)'s another pair that have a composition equal to \(h(x)\) are; \[g(x)=x^2\qquad\qquad f(x)=x^{3/2}+1\] \[f \circ g=f(g(x))=f(x^2)=(x^2)^{3/2}+1=x^3+1=h(x)\]

OpenStudy (unklerhaukus):

i am not sure how to find the general solution

OpenStudy (anonymous):

Not necessary. Im only required to enter one of the many solutions, the one you gave me works fine. Its the answer i got by myself but i had gotten the functions mixed. Thanx again

OpenStudy (unklerhaukus):

cool,